A room has AC which runs for 5hours a day at a voltage of 220V. The wiring of room consists of CU of radius 1mm and a length of 10m. Power consumption per day is 10 commercial units. What fraction it goes to joule heating in whires??what would happen if hte wiring of Al of the same dimensions are used??

Grade:12

4 Answers

Pushkar Aditya Test tutors
9 years ago
fdgd
Pushkar Aditya Test tutors
9 years ago
fdgd
Pushkar Aditya Test tutors
9 years ago
Tushar Nandanwar
5 years ago
Power consumption =10 units5hour=10 units 1 hour=2kW=2000J/sSo I=P/V=2000/220=9.09 AA Power loss in copper wire=I*I*R=I*Iρl/APower loss in copper wire=9.09×9.09×1.7×10−8×10π×(10−3)2=4.47J or 4.472000×100%=0.22% Power loss in aluminium wire=I*IR=I*Iρl/APower loss in copper wire=9.09×9.09×2.7×10−8×10π×(10−3)2=7.1J or 7.12000×100%=0.355%